A physical examination was performed for 8814 adult persons who visited the City Obesity Center during the month of June. Among these adults, 2089 had metabolic syndrome. Of these 2089 with metabolic syndrome, 1149 were obese. Of those adults without metabolic syndrome, 2286 were obese.

Answer these questions:

a) what is the study design?

b) generate a table (at least 2 by 2),

c) what is the prevalence of metabolic syndrome and what is prevalence of obesity in all visitors?

d) calculate prevalence odds ratios of obesity in subjects with metabolic syndrome compared to those without,

e) interpret your result in one sentence.
Answer #1

b)

had metabolic syndrome did not have metabolic syndrome
obese 1149 2286
not obese 940 4439
2089 6725

c)

prevalence of metabolic syndrome = 1149/2089 = 0.55

prevalence of obesity in all visitors = (1149 + 2286)/8814

= 0.3897

d)

odds ratio = 0.3897/(2286/6725) = 1.1464

e)

those with metabolic syndrome has 1.1464 times chance of being obese than that of without metabolic syndrome

Please rate